If a certain number x is subtracted from 8, it will be equal to two less than four times the same number. The correct equation for the above sentence is:

1
Expert's answer
2022-02-22T13:02:43-0500

If a certain number x is subtracted from 8, it will be equal to two less than four times the same number, or:

8 - x = 2 - 4x

or

x - 8 = 4x - 2.